# EchoTrail-GUI: Building Actionable Memory for GUI Agents via Critic-Guided Self-Exploration

### v1

@@ 1 Introduction
The rapid advancement of Large Vision-Language Models (VLMs) [bai2025qwen2, wang2024cogvlm] has fundamentally reshaped the development of autonomous GUI agents. Early approaches relied on structured metadata such as accessibility trees [wen2024autodroid], whereas modern systems can now parse raw GUI screenshots and follow complex user instructions through sequences of fine-grained actions, including clicking, scrolling, typing, and navigation [hong2024cogagent, qin2025ui, xu2024aguvis]. These capabilities have enabled significant progress in digital task automation, intelligent computer-use assistants, and interactive multimodal copilots.
@@
The first is the experience acquisition bottleneck. High-quality trajectory data is crucial for building generalizable GUI agents [rawles2024androidworld, xu2024androidlab], yet obtaining such data is expensive. Manual annotation is time-consuming and unscalable, while unguided exploration often produces incoherent or low-utility interaction traces. Recent efforts attempt to synthesize trajectories from videos [jang2025scalable], web tutorials [xu2024agenttrek, zhang2025tongui], or reverse task inference [sun2025genesis]. Others explore autonomous exploration [xie2025gui], but typically lack effective mechanisms to control trajectory quality or prevent the accumulation of noisy behaviors.
The second challenge is the knowledge application gap. Even when trajectory corpora are available, efficiently leveraging them remains nontrivial. Although VLMs possess strong in-context learning abilities, most GUI agents rely on static examples or handcrafted prompts [yang2023set, yang2025aria], failing to retrieve and apply past experience dynamically. Retrieval-augmented methods [xu2025retrieval] offer a promising direction, yet their effectiveness hinges on the availability of well-curated, semantically aligned operational memories—resources that are typically scarce or costly to obtain.
To address these challenges, we introduce EchoTrail-GUI, a unified framework that emulates the human-like cognitive cycle of learn–remember–apply. Our central goal is to construct a self-improving, closed-loop system where the agent autonomously gathers experience, distills actionable knowledge, and reuses it to guide new tasks. Inspired by retrieval-augmented generation (RAG) [lewis2020retrieval], EchoTrail-GUI enables GUI agents to dynamically access the most relevant and successful past trajectories as structured, contextual guidance.
@@
  • Through extensive experiments on AndroidWorld [rawles2024androidworld] and AndroidLab [xu2024androidlab], we demonstrate that memory-augmented GUI agents significantly outperform their stateless counterparts, yielding substantial gains in success rate, sub-goal completion, and operational robustness.
@@ 2 Related Work
GUI agents are designed to perceive and interact with digital environments by processing visual screen information. The emergence of powerful Multimodal Large Language Models (MLLMs) [achiam2023gpt, bai2025qwen2, anthropic2024Claude] has fundamentally shifted the paradigm from relying on structured metadata (like accessibility trees or XML) to end-to-end approaches that directly process raw screen pixels [rawles2024androidworld, wanyan2025look, lai2025androidgen]. Research in this area is primarily explored along two vectors. One major research thrust focuses on enhancing the intrinsic capabilities of a single powerful agent. This involves extensive pre-training on large, GUI-centric datasets [qin2025ui] or proposing unified action spaces to improve interaction abilities [wu2024atlas]. A parallel, training-free strategy leverages the in-context learning and reasoning of MLLMs, often combining reasoning approaches like ReAct [yao2022react] with visual annotation techniques [rawles2024androidworld, yang2023set]. The second vector explores the efficacy of collaborative multi-agent systems. These frameworks typically employ an orchestrator-agent architecture to decompose complex tasks and delegate sub-tasks [fourney2024magentic, lai2025androidgen]. In the GUI domain, this has been extended to integrating specialized modules such as reflection agents for performance evaluation and error correction [wang2024mobile], or self-evolving modules that utilize web knowledge and episodic memory for adaptation [zhang2025appagent].
@@
The creation of high-fidelity trajectory data is a critical prerequisite for advancing GUI agents, as this data underpins their ability to plan and execute tasks in complex digital worlds [pan2024autonomous]. The challenges associated with manual annotation have accelerated research into automated data generation. One approach involves synthesizing trajectories from existing human-generated content. This includes extracting (frame, action) sequences from online videos [jang2025scalable] or converting web tutorials into structured, synthetic trajectories suitable for agent training [xu2024agenttrek, zhang2025tongui]. A distinct method empowers agents to generate trajectories by autonomously interacting with the GUI environment. This ranges from simple random or seeded exploration [zhang2025appagent, wen2024autodroid] to more sophisticated techniques like ”reverse task synthesis,” which infers meaningful high-level tasks from observed low-level interaction traces [sun2025genesis].
@@
Moving beyond using external content solely for training, another paradigm treats these resources as dynamic, non-parametric sources of knowledge accessible at runtime. This Retrieval-Augmented Generation (RAG) approach aims to enhance agent performance without necessitating model fine-tuning. For example, some work employs lightweight models to retrieve relevant guides from a knowledge base, which are then summarized to instruct a frozen agent [xu2025retrieval]. Other research focuses on mining ”transition-aware knowledge” from autonomous exploration traces and formatting this knowledge into goal-oriented prompts to guide the agent’s decision-making during execution [xie2025gui].
@@ 4.1.1 Environments and Baselines
We conduct our evaluation on two widely-adopted, interactive Android benchmarks: AndroidWorld [rawles2024androidworld], which features a diverse set of real-world applications, and AndroidLab [xu2024androidlab], providing a more controlled environment for systematic testing.
Our framework’s performance is compared against a comprehensive suite of recent and state-of-the-art agents. This includes powerful proprietary models like GPT-4o and Gemini-1.5-Pro, as well as specialized open-source agents such as UGround [gou2024navigating] and UI-TARS [qin2025ui]. To demonstrate the model-agnostic nature of EchoTrail-GUI, we implement our framework on two distinct backbones: a leading open-source model, Qwen2.5-VL-72B-Instruct, and a leading proprietary model, GPT-4o.
@@ 4.2.2 Performance on AndroidLab
Closed Source AppAgent [zhang2025appagent] GPT-4o 14.9
Gemini [team2024gemini] Gemini-1.5-Pro 22.8
Claude [anthropic2024Claude] Claude Computer-Use 27.9
UGround [gou2024navigating] GPT-4o + UGround 32.8
GPT-4o [achiam2023gpt] GPT-4o 34.5
Ponder&Press [wang2025ponder] GPT-4o + Ponder&Press 34.5
Aguvis [xu2024aguvis] GPT-4o + Aguvis 37.1
M3A [rawles2024androidworld] GPT-4o 40.5
ScaleTrack [huang2025scaletrack] GPT-4o + ScaleTrack 44.0
Aria-UI [yang2025aria] GPT-4o + Aria-UI 44.8
URST [chenenhancingurst] GPT-4o + URST Reflexion 46.6
AndroidGen [lai2025androidgen] GPT-4o 46.8
GUI-explorer [xie2025gui] GPT-4o 47.4
@@
Open Source Aguvis [xu2024aguvis] Aguvis-72B 26.1
Qwen2.5-VL [bai2025qwen2] Qwen2.5-VL-72B-Instruct 35.0
RAG-GUI [xu2025retrieval] RAG-GUI-72B-RSF 45.7
UI-TARS [qin2025ui] UI-TARS-72B-SFT 46.6
@@
Closed Source Gemini[team2023gemini] Gemini-1.0 12.6 72.5 76.7 10.9
Claude [anthropic2024Claude] Claude-3-Opus 15.1 81.4 83.9 13.0
Gemini[team2024gemini] Gemini-1.5-Pro 18.5 106.0 91.5 16.7
Claude [anthropic2024Claude] Claude-3.5-Sonnet 32.7 113.4 81.2 29.0
GPT-4o [achiam2023gpt] GPT-4o 35.0 87.3 85.4 31.2
AutoGLM [liu2024autoglm] AutoGLM 36.2
@@
Open Source UI-TARS [qin2025ui] UI-TARS-72B 12.0 40.8 70.0 9.6
UI-TARS [qin2025ui] UI-TARS-72B-ft 28.4 81.4 81.6 22.1
Qwen2.5-VL [bai2025qwen2] Qwen2.5-VL-72B-Instruct 26.1 68.7 81.4 23.9
Qwen2-VL[wang2024qwen2] Qwen2-VL-72B-Instruct-ft 29.3 84.5 90.2 24.6
Qwen2.5-VL [bai2025qwen2] Qwen2.5-VL-72B-Instruct-ft 30.9 81.3 79.3 25.0
@@ 4.3.2 Analysis of the Self-Exploration Mechanism
A key claim of our work is that critic-guided self-exploration can autonomously generate a high-quality and diverse memory base. To validate this, we analyze the semantic properties of the trajectories in our generated EchoTrail-4K dataset. We use Sentence-BERT [reimers2019sentence] to encode the final task intents from our explored trajectories and compare them against the ground-truth task instructions from the AndroidLab benchmark. The resulting embeddings are then visualized in two dimensions using UMAP [mcinnes2018umap], as shown in Figure 2.

### v2

@@ 1 Introduction
The rapid advancement of Large Vision-Language Models (VLMs) [3, 23] has fundamentally reshaped the development of autonomous GUI agents. Early approaches relied on structured metadata such as accessibility trees [26], whereas modern systems can now parse raw GUI screenshots and follow complex user instructions through sequences of fine-grained actions, including clicking, scrolling, typing, and navigation [7, 15, 32]. These capabilities have enabled significant progress in digital task automation, intelligent computer-use assistants, and interactive multimodal copilots.
@@
The first is the experience acquisition bottleneck. High-quality trajectory data is crucial for building generalizable GUI agents [16, 30], yet obtaining such data is expensive. Manual annotation is time-consuming and unscalable, while unguided exploration often produces incoherent or low-utility interaction traces. Recent efforts attempt to synthesize trajectories from videos [9], web tutorials [31, 36], or reverse task inference [18]. Others explore autonomous exploration [28], but typically lack effective mechanisms to control trajectory quality or prevent the accumulation of noisy behaviors.
The second challenge is the knowledge application gap. Even when trajectory corpora are available, efficiently leveraging them remains nontrivial. Although VLMs possess strong in-context learning abilities, most GUI agents rely on static examples or handcrafted prompts [33, 34], failing to retrieve and apply past experience dynamically. Retrieval-augmented methods [29] offer a promising direction, yet their effectiveness hinges on the availability of well-curated, semantically aligned operational memories—resources that are typically scarce or costly to obtain.
To address these challenges, we introduce EchoTrail-GUI, a unified framework that emulates the human-like cognitive cycle of learn–remember–apply. Our central goal is to construct a self-improving, closed-loop system where the agent autonomously gathers experience, distills actionable knowledge, and reuses it to guide new tasks. Inspired by retrieval-augmented generation (RAG) [11], EchoTrail-GUI enables GUI agents to dynamically access the most relevant and successful past trajectories as structured, contextual guidance.
@@
  • Through extensive experiments on AndroidWorld [16] and AndroidLab [30], we demonstrate that memory-augmented GUI agents significantly outperform their stateless counterparts, yielding substantial gains in success rate, sub-goal completion, and operational robustness.
@@ 2 Related Work
GUI agents are designed to perceive and interact with digital environments by processing visual screen information. The emergence of powerful Multimodal Large Language Models (MLLMs) [1, 3, 2] has fundamentally shifted the paradigm from relying on structured metadata (like accessibility trees or XML) to end-to-end approaches that directly process raw screen pixels [16, 25, 10]. Research in this area is primarily explored along two vectors. One major research thrust focuses on enhancing the intrinsic capabilities of a single powerful agent. This involves extensive pre-training on large, GUI-centric datasets [15] or proposing unified action spaces to improve interaction abilities [27]. A parallel, training-free strategy leverages the in-context learning and reasoning of MLLMs, often combining reasoning approaches like ReAct [35] with visual annotation techniques [16, 33]. The second vector explores the efficacy of collaborative multi-agent systems. These frameworks typically employ an orchestrator-agent architecture to decompose complex tasks and delegate sub-tasks [5, 10]. In the GUI domain, this has been extended to integrating specialized modules such as reflection agents for performance evaluation and error correction [21], or self-evolving modules that utilize web knowledge and episodic memory for adaptation [37].
@@
The creation of high-fidelity trajectory data is a critical prerequisite for advancing GUI agents, as this data underpins their ability to plan and execute tasks in complex digital worlds [14]. The challenges associated with manual annotation have accelerated research into automated data generation. One approach involves synthesizing trajectories from existing human-generated content. This includes extracting (frame, action) sequences from online videos [9] or converting web tutorials into structured, synthetic trajectories suitable for agent training [31, 36]. A distinct method empowers agents to generate trajectories by autonomously interacting with the GUI environment. This ranges from simple random or seeded exploration [37, 26] to more sophisticated techniques like ”reverse task synthesis,” which infers meaningful high-level tasks from observed low-level interaction traces [18].
@@
Moving beyond using external content solely for training, another paradigm treats these resources as dynamic, non-parametric sources of knowledge accessible at runtime. This Retrieval-Augmented Generation (RAG) approach aims to enhance agent performance without necessitating model fine-tuning. For example, some work employs lightweight models to retrieve relevant guides from a knowledge base, which are then summarized to instruct a frozen agent [29]. Other research focuses on mining ”transition-aware knowledge” from autonomous exploration traces and formatting this knowledge into goal-oriented prompts to guide the agent’s decision-making during execution [28].
@@ 4.1.1 Environments and Baselines
We conduct our evaluation on two widely-adopted, interactive Android benchmarks: AndroidWorld [16], which features a diverse set of real-world applications, and AndroidLab [30], providing a more controlled environment for systematic testing.
Our framework’s performance is compared against a comprehensive suite of recent and state-of-the-art agents. This includes powerful proprietary models like GPT-4o and Gemini-1.5-Pro, as well as specialized open-source agents such as UGround [6] and UI-TARS [15]. To demonstrate the model-agnostic nature of EchoTrail-GUI, we implement our framework on two distinct backbones: a leading open-source model, Qwen2.5-VL-72B-Instruct, and a leading proprietary model, GPT-4o.
@@ 4.2.2 Performance on AndroidLab
Closed Source AppAgent [37] GPT-4o 14.9
Gemini [20] Gemini-1.5-Pro 22.8
Claude [2] Claude Computer-Use 27.9
UGround [6] GPT-4o + UGround 32.8
GPT-4o [1] GPT-4o 34.5
Ponder&Press [24] GPT-4o + Ponder&Press 34.5
Aguvis [32] GPT-4o + Aguvis 37.1
M3A [16] GPT-4o 40.5
ScaleTrack [8] GPT-4o + ScaleTrack 44.0
Aria-UI [34] GPT-4o + Aria-UI 44.8
URST [4] GPT-4o + URST Reflexion 46.6
AndroidGen [10] GPT-4o 46.8
GUI-explorer [28] GPT-4o 47.4
@@
Open Source Aguvis [32] Aguvis-72B 26.1
Qwen2.5-VL [3] Qwen2.5-VL-72B-Instruct 35.0
RAG-GUI [29] RAG-GUI-72B-RSF 45.7
UI-TARS [15] UI-TARS-72B-SFT 46.6
@@
Closed Source Gemini[19] Gemini-1.0 12.6 72.5 76.7 10.9
Claude [2] Claude-3-Opus 15.1 81.4 83.9 13.0
Gemini[20] Gemini-1.5-Pro 18.5 106.0 91.5 16.7
Claude [2] Claude-3.5-Sonnet 32.7 113.4 81.2 29.0
GPT-4o [1] GPT-4o 35.0 87.3 85.4 31.2
AutoGLM [12] AutoGLM 36.2
@@
Open Source UI-TARS [15] UI-TARS-72B 12.0 40.8 70.0 9.6
UI-TARS [15] UI-TARS-72B-ft 28.4 81.4 81.6 22.1
Qwen2.5-VL [3] Qwen2.5-VL-72B-Instruct 26.1 68.7 81.4 23.9
Qwen2-VL[22] Qwen2-VL-72B-Instruct-ft 29.3 84.5 90.2 24.6
Qwen2.5-VL [3] Qwen2.5-VL-72B-Instruct-ft 30.9 81.3 79.3 25.0
@@ 4.3.2 Analysis of the Self-Exploration Mechanism
A key claim of our work is that critic-guided self-exploration can autonomously generate a high-quality and diverse memory base. To validate this, we analyze the semantic properties of the trajectories in our generated EchoTrail-4K dataset. We use Sentence-BERT [17] to encode the final task intents from our explored trajectories and compare them against the ground-truth task instructions from the AndroidLab benchmark. The resulting embeddings are then visualized in two dimensions using UMAP [13], as shown in Figure 2.